### v1

# S-GRPO: Early Exit via Reinforcement Learning in Reasoning Models

Abstract
As Test-Time Scaling emerges as an active research focus in the large language model community, advanced post-training methods increasingly emphasize extending chain-of-thought (CoT) generation length, thereby enhancing reasoning capabilities to approach Deepseek R1-like reasoning models. However, recent studies reveal that reasoning models (even Qwen3) consistently exhibit excessive thought redundancy in CoT generation. This overthinking problem stems from conventional outcome-reward reinforcement learning’s systematic neglect in regulating intermediate reasoning steps. This paper proposes Serial-Group Decaying-Reward Policy Optimization (namely S-GRPO), a novel reinforcement learning method that empowers models with the capability to determine the sufficiency of reasoning steps, subsequently triggering early exit of CoT generation. Specifically, unlike GRPO, which samples multiple possible completions (parallel group) in parallel, we select multiple temporal positions in the generation of one CoT to allow the model to exit thinking and instead generate answers (serial group), respectively. For the correct answers in a serial group, we assign rewards that decay according to positions, with lower rewards towards the later ones, thereby reinforcing the model’s behavior to generate higher-quality answers at earlier phases with earlier exits of thinking. Empirical evaluations demonstrate compatibility with state-of-the-art reasoning models, including Qwen3 and Deepseek-distill models, achieving 35.4$\sim$61.1sequence length reduction with 0.72$\sim$6.08accuracy improvements across GSM8K, AIME 2024, AMC 2023, MATH-500, and GPQA Diamond benchmarks.
Figure 1: Comparison of parallel-group-relative GRPO and our serial-group-relative S-GRPO.
1 Introduction
Test-Time Scaling snell2024scalingllmtesttimecompute demonstrates a strong correlation between extended chain-of-thought (CoT) and enhanced reasoning capabilities in Large Language Models, which is more effective than scaling model parameters kaplan2020scalinglawsneurallanguage . The success of DeepSeek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability and GPT-o1 openai2025learning has further sparked the research interest in reasoning models xu2025largereasoningmodelssurvey within the LLM community. Furthermore, reinforcement learning shao2024deepseekmathpushinglimitsmathematical ; bai2022training ; ouyang2022training ; schulman2017proximal ; ramesh2024group (RL) in post-training has demonstrated significant potential in stimulating long chain-of-thought generation and strengthening deep thinking.
However, recent studies chen2025think23overthinkingo1like ; team2025kimi have uncovered a critical limitation in reasoning models: their inherent tendency toward redundant thought processes, namely Overthinking chen2025think23overthinkingo1like ; cuadron2025danger . They frequently generate unnecessarily lengthy reasoning sequences wu2025more ; cuadron2025danger that incorporate extraneous information and non-essential reasoning steps. This redundant thought not only inflates computational overhead but also reduces reasoning accuracy through deviation from valid reasoning pathways to incorrect conclusions yang2025dynamicearlyexitreasoning . We attribute this phenomenon to the inherent nature of 0/1 outcome-reward RL (exemplified by GRPO shao2024deepseekmathpushinglimitsmathematical ), where training with only final outcome rewards is unable to determine when reasoning steps have reached sufficiency.
As shown in Figure 1 (left), the standard outcome-reward RL, like GRPO, tends to sample a query multiple times in parallel to obtain a parallel group. All reasoning chains and corresponding conclusions within the parallel group receive 0/1 outcome rewards, which reinforce correct ones regardless of overthinking in intermediate steps. While this approach successfully aggregates pass@k reasoning capability into pass@1, the neglect of regulating intermediate reasoning results in inefficient inference. Conversely, as illustrated in Figure 1 (right), we constructe a serial group by sequentially collecting multiple completions during single CoT generation, and prioritize rewards for earlier correct ones, aiming to enable models to acquire and exert the complete reasoning capability in the initial generation phases, triggering early exit before overthinking occurs.
Motivated by this, we propose Serial-Group Decaying-Reward Policy Optimization (S-GRPO), a simple yet effective modification to address standard outcome-reward RL (GRPO)’s inability to regulate intermediate reasoning processes. Specifically, during S-GRPO training, we construct a serial group for each query through a two-phase rollout: the initial rollout generates a complete reasoning path, followed by subsequent rollouts that perform early-exit interventions at different positions to generate intermediate answers, which are collected with their corresponding truncated reasoning paths into the serial group. On this basis, we assign the rewards that decay according to their order of early exit for the correct ones within the serial group. S-GRPO culminates in computing serial-group relative advantages and using their policy gradient to update model parameters.
Our contributions are threefold:
  • We pioneer a serial-group RL paradigm that overcomes the critical limitation of outcome-reward RL in regulating intermediate reasoning processes, accompanied by an open-sourced training framework.
  • The proposed S-GRPO algorithm empowers models to generate higher-quality reasoning paths at early stages of CoTs while implementing early exit when the sufficiency of generated thought is met – a strategy particularly suitable as final-stage optimization in post-training pipelines.
  • Extensive experiments across MATH-500, AMC, GSM8K, AIME and GPQA benchmarks with Qwen3 and Deepseek-series reasoning models, demonstrate 0.74$\sim$6.08accuracy improvement alongside 35.4$\sim$61.1average token reduction, establishing an efficiency-accuracy synergistic improvement.
2 Method
[…]
The proposed Serial-Group Decaying-Reward Policy Optimization (S-GRPO) is a novel reinforcement learning mechanism that innovatively leverages rule-based outcome rewards to regulate intermediate reasoning processes. During the training of S-GRPO, we exit early and generate answers at different positions in a complete CoT, thus constructing a serial group and assigning rewards that decrease according to the position, which reinforces the model to generate high-quality thoughts earlier and exit early when the generated thought is sufficient. In particular, the framework is divided into three stages: Serial-Group Generation, Decaying Reward Policy, and Advantage Computation and Parameter Update.
2.1 Serial-Group Generation
GRPO is originally designed to enable reasoning models to achieve their pass@k potentiality under pass@1 settings. To accomplish this, GRPO performs parallel sampling of multiple CoTs per query to construct a parallel group, then rewards the correct ones. In contrast, S-GRPO aims to achieve efficient reasoning through sufficient-reasoning early exit. Specifically, we perform early-exit interventions at different positions within a single CoT to construct a serial group, thereby allowing the RL training can compare the thought sufficiency between reasoning paths of different lengths. It consists of two stages: Full Thought Rollout and Early-exit Thought Rollout.
2.1.1 Full Thought Rollout
In the Full Thought Rollout, the model generates a complete reasoning path step by step $\{O^{0}=T_{1},T_{2},\ldots,T_{n},</think>,C_{0}\}$ for each query. To simulate diverse reasoning sufficiency scenarios and train the model to handle early exits effectively, we adopt random-length truncation during training. Specifically, the reasoning sequence is randomly truncated at $m$ different temporal positions $P_{i}$, where $P_{i}=T_{i}\quad and\quad i\sim\text{Uniform}(1,n).$ This randomness ensures that the model learns to evaluate sufficiency implicitly across varying reasoning depths and avoids overfitting to fixed-length reasoning paths.
2.1.2 Early-exit Thought Rollout
In the Early-exit Thought Rollout, the policy model extracts several early-exit thougths $\{CoT^{1},CoT^{2},...,CoT^{m}\}=\{O^{0}[:P_{1}],O^{0}[:P_{2}]...,,O^{0}[:P_{m}]\}$ of different lengths between the first token and multiple trancated-postion tokens of the full thought, respectively, and generates answers $C_{1},C_{2},\ldots,C_{m}$ for all early-exit CoTs.
[…]
To ensure that the serial group contains early exit samples with correct answers (obtained by Early-exit Thought Rollout), we sample more queries than the training batch required (over-sampling) for data filtering, such as DAPO yu2025dapo .
2.2 Decaying Reward Policy
To encourage models to produce adequate and correct reasoning steps at earlier stages of CoT generation for accurate problem-solving, we propose a Decaying Reward Policy. This mechanism assigns rewards based on the correctness of answers generated from the two-time rollouts ( $C_{1},C_{2},\ldots,C_{m},and\ C_{0}$ ), while decaying rewards according to their order of early exits. For the answer $C^{i}$ of each response $O^{i}$, the reward $r^{i}$ is defined as follows:
$r^{i}=\begin{cases}\frac{1}{2^{N_{\text{right}}-1}},&\text{if }C^{i}\text{ is correct,}\\ 0,&\text{if }C^{i}\text{ is incorrect.}\end{cases}$
Where $N_{\text{right}}$ refers to the accumulated number of correct answers up to and including the current position. The Decaying Reward Policy is designed with dual objectives: (1) Exponential decay for correct answers: We employ exponentially diminishing rewards to prioritize the quality of earlier reasoning steps, which is overlooked by binary 0/1 outcome rewards. (2) Zero reward for incorrect answers: This ensures correctness-first optimization, preventing degradation of the model’s core reasoning capabilities.
The design of the above two stages balances the trade-off between reasoning sufficiency and efficiency, guiding the model toward producing accurate and concise reasoning sequences.
2.3 Advantage Computation and Parameter Update
After obtaining the decaying rewards, GRPO calculates the advantage for each sample based on the serial group’s rewards. Specifically, the mean of the rewards within the group is computed, and the advantage for each sample is calculated using the formula: $\hat{A}_{i}=r_{i}-\text{mean}(r_{i})$. Here, for training stability, the standard deviation is removed from the advantage computation compared to the GRPO. Subsequently, the computed advantage for each sample is broadcast to all corresponding response tokens. Finally, parameter updates are performed based on the advantage values of each sample (Algorithm 1). The optimization objective is as follows:
$\mathcal{J}_{\text{S-GRPO}}(\theta)=\mathbb{E}_{[q\sim P(Q),\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(O\mid q)]}\\ [\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\{\min[\frac{\pi_{\theta}^{i,t}}{\pi_{\theta_{\text{old}}}^{i,t}}\hat{A}_{i,t},\,\text{clip}(\frac{\pi_{\theta}^{i,t}}{\pi_{\theta_{\text{old}}}^{i,t}},\,1-\epsilon,\,1+\epsilon)\hat{A}_{i,t}]\}]$ (1)
where $\pi^{i,t}=\pi(o_{i,t}\mid q,o_{i,<t})$ denotes the conditional probability of the token at position $t$. The models $\pi_{\theta_{d}}$, $\pi_{{\theta_{d}}_{\mathrm{old}}}$ correspond to the training model and sampling model, respectively. $q$ represents the input query, and $\{o_{i}\}_{i=1}^{G}$ are the full and early-exit thought rollout responses generated by the model. The advantage $\hat{A}_{i}$ for each response is computed as $\hat{A}_{i}=r_{i}-\text{mean}(r_{i})$, where $r_{i}$ is the decaying reward assigned to the response. For token-level advantage $\hat{A}_{i,t}$, it is defined to be equal to the corresponding sequence-level advantage $\hat{A}_{i}$. The hyperparameter $\epsilon$ is used to bound the importance sampling ratio.
Algorithm 1 Serial-Group Decaying-Reward Policy Optimization (S-GRPO)
1:Query $Q$, Policy model $\pi_{\theta}$, Number of positions to sample $m$
2:Updated policy parameters $\theta$
3:Full Thought Rollout:
4:Generate complete reasoning path $O^{0}=(T_{1},T_{2},\ldots,T_{n},\text{</think>},C^{0})$ using $\pi_{\theta}$ for query $Q$
5:Sample $m$ temporal positions $\{P_{1},P_{2},\ldots,P_{m}\}$ where $P_{i}=T_{i}$ and $i\sim\text{Uniform}(1,n)$
6:Early-exit Thought Rollout:
[…]
8:for each position $P_{i}$ where $i\in\{1,2,\ldots,m\}$ do
9:     Append $T_{\text{</think>}}$ after position $P_{i}$ to form $CoT^{i}=(T_{1},T_{2},\ldots,T_{i},T_{\text{</think>}})$
10:     Generate answer $C^{i}$ using policy model $\pi_{\theta}$ conditioned on $CoT^{i}$
11:     Form output $O^{i}=(CoT^{i},C^{i})$
12:end for
13:Rule-based Dynamic Decaying-reward:
14:$N_{\text{right}}\leftarrow 0$triangleright Counter for correct answers
15:for each output $O^{i}$ where $i\in\{1,2,\ldots,m,0\}$ in order of position do
16:     if $C^{i}$ is correct then
17:         $N_{\text{right}}\leftarrow N_{\text{right}}+1$
18:         $r^{i}\leftarrow\frac{1}{2^{N_{\text{right}}-1}}$triangleright Exponentially decaying reward
19:     else
20:         $r^{i}\leftarrow 0$triangleright Zero reward for incorrect answers
21:     end if
22:end for
23:Group Computation:
24:Compute advantages $A^{i}$ for each output based on rewards $r^{i}$
25:Update policy parameters $\theta$ using policy gradient with advantages $A^{i}$
3 Experiments
3.1 Experimental Setup
Training datasets.
We selected problems from DeepMath-103K he2025deepmath103klargescalechallengingdecontaminated to build our training set. This dataset is a large-scale and challenging collection of mathematics problems, focusing on difficulty levels ranging from grade 5 to grade 10. It addresses the lack of sufficient complexity present in existing datasets by offering approximately 103,000 carefully curated problems. The dataset is constructed through an extensive data cleaning process applied to known benchmarks, ensuring no overlap with widely used benchmarks.
Benchmarks.
To comprehensively assess the models’ capabilities across a range of reasoning tasks, we select five popular benchmarks that reflect diverse levels of difficulty: GSM8K (cobbe2021trainingverifierssolvemathgsm8k, ) is a well-curated collection of 1,319 problems in elementary mathematics. This benchmark is specifically designed to evaluate multi-step reasoning in foundational math tasks. Problems typically require between two and eight sequential operations, relying primarily on basic arithmetic performed over multiple intermediate steps. AIME 2024 (aime, ) includes 30 advanced problems selected from the 2024 edition of the American Invitational Mathematics Examination. This highly regarded competition tests participants’ ability to reason mathematically across a broad range of topics, including arithmetic, algebra, combinatorics, geometry, number theory, and probability—core components of secondary-level mathematics. AMC 2023 (AMC2023, ) consists of 40 questions spanning key areas such as algebra, geometry, number theory, and combinatorics. As part of the American Mathematics Competitions organized by the Mathematical Association of America (MAA), AMC aims to foster problem-solving skills and identify promising young mathematicians. MATH-500 math500hendrycks2021measuringmathematicalproblemsolving presents a set of challenging problems drawn from high school-level mathematical competitions. To ensure comparability with prior studies, we use the same subset of 500 problems originally compiled by OpenAI for evaluation purposes (lightman2023letsverifystepstep, ). In addition to these mathematical evaluations, we also examine performance on scientific reasoning tasks, using the following benchmark: GPQA (rein2023gpqagraduatelevelgoogleproofqa, ) is a rigorously constructed dataset containing graduate-level questions in physics, chemistry, and biology. Notably, even domain experts with PhDs achieve only 69.7% accuracy on this benchmark o1 . For our experiments, we focus on the highest-quality subset, known as GPQA Diamond, which contains 198 carefully chosen questions.
Baselines.
We compare S-GRPO with various existing efficient reasoning methods, including the training-free, output-based approach DEER yang2025dynamicearlyexitreasoning , the off-policy preference optimization method ConCISE qiao2025conciseconfidenceguidedcompressionstepbystep , and on-policy RL-based approaches such as original GRPO, RL + Length Penalty arora2025traininglanguagemodelsreason , and ShorterBetter yi2025shorterbetterguidingreasoningmodels . DEER makes early-exit decisions during the inference phase, guided by the confidence scores of intermediate answers. ConCISE generates more concise chains of thought by integrating specialized prompt tokens and applying early-exit mechanisms during inference, followed by preference optimization using SimPO to further encourage succinct reasoning. RL + Length Penalty assigns reward values based on the deviation of each individual response length from the mean, penalizing longer responses. ShorterBetter promotes shorter yet accurate reasoning paths by assigning higher rewards to compact chains that yield correct answers based on GRPO.
Models.
We evaluate S-GRPO on four large reasoning models, including DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-14B deepseekai2025deepseekr1incentivizingreasoningcapability , Qwen3-8B, and Qwen3-14B qwen2025qwen25technicalreport . Despite achieving state-of-the-art performance on reasoning tasks, these models tend to generate overly verbose reasoning processes that contain excessive redundant information.
Metrics.
S-GRPO aims to enhance the correctness performance while reducing inference length, thereby enabling more efficient reasoning. To this end, we selected Accuracy (i.e., pass@1) and Token Number (Tokens) as the evaluation metrics. Given the instability associated with generating long sequences in large reasoning models, as well as the limited number of samples in certain benchmark datasets, we conducted multiple runs for each evaluation and report the averaged results in the tables. Specifically, we ran 16 trials on AIME-2024 and AMC-2023, 8 trials on MATH-500 and GPQA Diamond, and 4 trials on GSM8K.
Training details.
During the S-GRPO, we use a learning rate of $1\times 10^{-6}$ and randomly choose 8 temporal positions for each query. As we use on-policy mode, the generation batch size and train batch size are the same and equal to 128*8. For the GRPO, we also use a learning rate of $1\times 10^{-6}$ and set the generation batch size and train batch size to 128*8. For the “RL + Length Penalty”, we refer to the settings in its paper arora2025traininglanguagemodelsreason and choose the scalar parameter ($\alpha$) of 0.2. For all experiments, Adam kingma2014adam is used as the standard optimizer.
Table 1: Experimental results on four Large Reasoning Models.
Method GSM8K AIME24 AMC23 MATH-500 GPQA Overall
Acc Tokens Acc Tokens Acc Tokens Acc Tokens Acc Tokens Acc Tokens
DeepSeek-R1-Distill-Qwen-7B
Vanilla 92.4 1833 55.4 13232 77.2 9693 85.8 5590 50.1 15385 70.02 9147
DEER 88.8 917 53.3 10971 87.5 4142 91.8 2431 47.5 5280 73.78+3.76 4748-48.1%
$\textit{ConCISE}_{SFT}$ 92.9 832 52.1 9751 92.0 2244 50.0 5892
$\textit{ConCISE}_{SimPO}$ 92.1 715 48.3 7745 91.0 1946 48.0 4859
GRPO 93.2 1767 55.0 13451 87.5 9887 93.6 5317 50.7 15817 76.00+5.98 9248+1.1%
RL + Length Penalty 92.4 1062 51.9 7464 86.9 3540 92.2 2451 49.1 3984 74.50+4.48 3700-59.5%
ShorterBetter 53.3 5288 75.9 2580
S-GRPO 93.8 906 56.0 7377 87.5 3494 92.4 2252 50.8 3751 76.10+6.08 3556-61.1%
DeepSeek-R1-Distill-Qwen-14B
Vanilla 94.2 2129 64.4 11099 90.5 5527 93.5 3844 59.2 6034 80.36 5727
DEER 93.3 982 70.0 10335 90.0 4349 91.4 2753 57.1 4767 80.36+0.0 4637-19.0%
GRPO 95.3 2120 65.8 13504 91.9 6595 84.0 4471 58.9 7354 79.18-1.18 6809+18.9%
RL + Length Penalty 94.7 775 55.0 7950 88.1 3396 92.4 1993 56.0 4380 77.24-3.12 3699-34%
S-GRPO 96.2 724 64.4 6712 91.9 3352 93.6 2146 59.3 3334 81.08+0.72 3253-35.4%
Qwen3-8B
Vanilla 95.4 2370 74.1 15326 91.3 9452 93.4 5577 55.6 8741 81.90 8293
DEER 95.5 981 76.7 11287 95.0 6198 93.4 3208 52.5 3104 82.62+0.72 4956-40.2%
GRPO 95.8 2355 72.7 15154 92.8 8983 94.4 5440 55.8 8819 82.30+0.4 8150-1.7%
RL + Length Penalty 95.4 1323 73.8 9666 93.4 5042 94.2 3247 56.2 5293 82.60+0.7 4914-40.7%
S-GRPO 96.1 1292 77.3 8810 95.0 5962 95.2 3166 57.7 5271 84.26+2.36 4922-40.6%
Qwen3-14B
Vanilla 95.5 1909 75.4 14116 96.9 7576 95.2 5078 58.8 7576 84.36 7251
DEER 95.5 908 76.7 10333 95.0 5099 94.8 2987 57.1 2435 83.82-0.54 4352-40.0%
GRPO 96.1 1956 77.7 14544 98.4 8000 95.8 5140 59.3 7966 85.46+1.1 7521+3.7%
RL + Length Penalty 95.8 1090 74.8 9056 96.6 5059 95.8 2866 59.4 4949 84.46+0.1 4604-36.5%
S-GRPO 96.3 952 77.9 8932 97.8 4537 96.4 2652 60.6 4537 85.80+1.14 4322-40.4%
3.2 Experimental Results
Main Results.
The experimental results in Table 2 demonstrate that our S-GRPO consistently and significantly outperforms existing baselines across five benchmark datasets and four reasoning models. Compared to vanilla reasoning models, S-GRPO achieves an average accuracy improvement of 0.74 to 6.08 percentage points, while reducing the generated sequence length by 35.4% to 61.1%. Our method delivers notable gains both on in-domain mathematical reasoning benchmarks used in training (i.e., GSM8K, AIME-2024, AMC-2023, MATH-500) and on out-of-domain scientific reasoning tasks (i.e., GPQA), providing strong evidence of its effectiveness and robustness. Specifically, when applied to DeepSeek-R1-Distill-Qwen-7B, S-GRPO achieves a 10.3 points increase in accuracy on AMC 2023 while utilizing only 36% of the original reasoning budget. Similarly, it obtains a 0.6 point improvement on AIME 2024 with just 56% of the allocated inference length.
Comparison with SoTAs.
In comparison, DEER reduces reasoning length by only 17% on AIME-2024, yet at the cost of a decline in accuracy. Unlike DEER, whose efficiency is primarily limited to simple tasks, S-GRPO consistently reduces reasoning length across both easy and challenging tasks, while also enhancing reasoning accuracy. Compared to the original GRPO, S-GRPO achieves comparable or even better accuracy while significantly shortening the inference trajectory, indicating that the proposed Serial-Group Generation mechanism does not hinder the model’s exploration capability in reinforcement learning. Moreover, the reward shaping based on serialized intermediate outputs better guides the model toward efficient reasoning.
Compared with the most recent training-based efficient reasoning methods, S-GRPO achieves the best performance on both reasoning length reduction and accuracy. This can be explained by the fact that ConCISE’s off-policy preference optimization forces the model to fit into a new data distribution, and ShorterBetter and RL + Length Penalty overly emphasize length reduction, at the expense of generalization and accuracy-driven reward optimization. In contrast, S-GRPO preserves the integrity of the original reasoning process via a two-stage rollout procedure, ensuring that the model’s pre-existing reasoning abilities are not compromised, which is well-suited as the final stage of post-training.
Figure 3: Performance of DeepSeek-R1-Distill-Qwen-7B and Qwen3-8B without or with S-GRPO training on GSM8K and AIME24 under different generation-length budgets.
Performance with different token budget.
We vary the generation-length budget during inference from short to long, and compare the accuracy and actual reasoning length of S-GRPO and vanilla CoT on GSM8K (representing simpler problems) and AIME-2024 (representing more complex problems). The results in Figure 3 show that, across all tested budgets, S-GRPO consistently outperforms vanilla CoT in both accuracy and efficiency with shorter generated sequences, further demonstrating the effectiveness of our method.
Moreover, we observe that under tight length budgets, S-GRPO achieves significantly higher accuracy than vanilla CoT while generating sequences of similar length. In contrast, under larger length budgets, S-GRPO produces significantly shorter reasoning paths with slightly better accuracy. The trends in both accuracy and actual generation length of S-GRPO are smoother compared to vanilla CoT, indicating its greater robustness to variations in length budget. Overall, S-GRPO achieves high accuracy with a much lower length budget, suggesting that our method is capable of generating concise yet accurate reasoning paths.
Table 2: Ablation results on Qwen3-8B.
Method GSM8K AIME24 AMC23 MATH-500 GPQA Overall
Acc Tokens Acc Tokens Acc Tokens Acc Tokens Acc Tokens Acc Tokens
Qwen3-8B
S-GRPO 96.1 1292 77.3 8810 95.0 5962 95.2 3166 57.7 5271 84.26 4922
w/o. Decaying (Shortest 1) 95.9 1175 69.6 8721 92.5 4581 94.8 2740 55.7 4734 81.70-2.56 4390-10.8%
w/o. Decaying (All 1) 96.0 2385 74.4 14940 94.7 9000 95.0 5614 54.9 8955 83.00-1.26 8179+66.2%
  - w/o. Serial 95.8 2355 72.7 15154 92.8 8983 94.4 5440 55.8 8819 82.30-1.96 8150+65.6%
Ablation Results.
To verify the effectiveness of each design component in S-GRPO, we conduct ablation studies under three different settings. w/o. Decaying (Shortest 1) denotes the setting where only the shortest correct completion sampled from the two-time rollouts is assigned a reward of 1, while all other samples receive a reward of 0. w/o. Decaying (All 1) refers to the configuration in which all correctly answered completions sampled via two-time rollouts are assigned a reward of 1, while incorrect ones receive a reward of 0. w/o. Serial indicates that, based on the w/o. Decaying (All 1) setup, we further remove the Serial-Group Generation mechanism.
The results in Table 2 indicate that assigning rewards only to the shortest correct completion imposes an overly strict constraint. Although this setting leads to further reductions in reasoning length, it comes at the cost of reasoning accuracy. Removing the design that assigns higher rewards to shorter outputs, i.e., w/o. Decaying (All 1), still results in lengthy reasoning. We attribute this to the fact that longer reasoning sequences are more likely to lead to correct answers, as suggested by test-time scaling theory, causing the model to ultimately converge toward exploring longer reasoning paths. When the Serial-Group Generation mechanism is removed, our method degenerates to GRPO, achieving performance comparable to w/o. Decaying (All 1) in both accuracy and reasoning length. This demonstrates that Serial-Group Generation, as an essential component of S-GRPO, does not compromise the model’s exploration capability in reinforcement learning.
Case Study.
Figure 5 in the Appendix visually illustrates how S-GRPO samples completions that exit early at different positions during the second rollout, along with the corresponding reward assignments. For the first early-exit position, where the model produces an incorrect intermediate answer, we set the reward to zero. For subsequent exits that yield correct answers, we apply a decaying positive reward scheme, where earlier exits are associated with higher reward values. This design incentivizes the model to discover reasoning paths that are both accurate and succinct.
To more intuitively illustrate the effectiveness of our approach, we present a representative example in Figure 4. The generated contents on the left and right sides of the figure illustrate the reasoning processes of S-GRPO and vanilla CoT, respectively. Although both methods yield the correct final answer, S-GRPO achieves this with less than half of the allocated reasoning budget, demonstrating its ability to effectively mitigate the overthinking problem chen2025think23overthinkingo1like . The central portion of the figure displays an early exit obtained by directly truncating the original model’s reasoning process at the same thinking budget as our method. The reasoning model fails to reach the correct conclusion based on the available partial reasoning. This indicates that S-GRPO more accurately identifies the correct solution path and guides the model toward concise yet accurate reasoning, inherently avoiding the underthinking issue where each reasoning path is only explored superficially wang2025thoughtsplaceunderthinkingo1like .
4 Related Work
The success of OpenAI o1 openai2025learning ; ouyang2022traininglanguagemodelsfollow highlights the powerful potential of reinforcement learning in post-training to enhance model reasoning capabilities. With the open source of Deepseek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability and Qwen3 qwen2025qwen25technicalreport , reasoning models are now widely deployed locally, drawing the LLM community’s attention to the efficiency of long chain-of-thought generation.
Reinforcement Learning for Reasoning
Rule-based outcome reward RL deepseekai2025deepseekr1incentivizingreasoningcapability ; kimiteam2025kimik15scalingreinforcement ; gao2024designingeffectiverlreward ; lambert2025tulu3pushingfrontiers ; zeng2025simplerlzooinvestigatingtamingzero ; wen2025lightr1curriculumsftdpo ; song2025fastcurlcurriculumreinforcementlearning has emerged as the mainstream approach for post-training optimization of reasoning models. This paradigm simplifies reward design by employing binary 0/1 rewards determined through rule-based correctness evaluation, eliminating the need for separate reward models as required in original GRPO shao2024deepseekmathpushinglimitsmathematical ; deepseekai2025deepseekr1incentivizingreasoningcapability ; schulman2017proximalpolicyoptimizationalgorithms implementations, thereby substantially reducing memory and computational overhead during RL training. Recent advances have introduced numerous RL algorithm variants focusing on training efficiency yu2025dapoopensourcellmreinforcement ; liu2025understandingr1zeroliketrainingcritical ; deepcoder2025 ; zhang2025srpocrossdomainimplementationlargescale , value function optimization kazemnejad2024vineppounlockingrlpotential ; yuan2025whatspposcollapselongcot ; yue2025vapoefficientreliablereinforcement , and other aspects. However, these developments preserve GRPO’s fundamental mechanism that samples several CoTs in parallel for a group, consequently overlooking optimization opportunities during sequential thought generation within a single CoT. Our proposed optimized version of GRPO, S-GRPO, maintains the rule-based outcome reward framework while enabling intermediate reasoning process rewards by sampling multiple early-exit thought chains in serial for one CoT. We allow existing reasoning models to effectively mitigate overthinking and boost inference efficiency.
Figure 4: Comparison of a generated content sample on GSM8K.
Efficient Reasoning
Approaches for efficient reasoning can be categorized into two groups: training-free and training-based methods. Training-free methods typically improve reasoning efficiency through dynamic prompting strategies han2024token ; xu2025chain ; lee2025well ; renze2024benefits ; chen2024unlocking ; aytes2025sketch ; chuang2024learning ; ong2406routellm , Best-of-N sampling pruning xie2023self ; liao2025reward and optimizations li2024escape ; manvi2024adaptive ; aggarwal2023let , and early-exit ma2025reasoning ; yang2025dynamicearlyexitreasoning mechanisms during reasoning. DEER yang2025dynamicearlyexitreasoning and our method employ a similar operation by implementing early exit during CoT generation and producing intermediate answers. The key distinction lies in that DEER directly makes early-exit decisions during inference based on the confidence of intermediate answers, whereas our method leverages RL to reinforce models’ behavior of correct early exit, which requires no additional compute overhead during inference.
Training-based methods primarily supervised fine-tuning models with variable-length CoT data yu2024distilling ; kang2025c3ot ; xia2025tokenskip ; ma2025cot ; munkhbat2025self ; liu2024can ; han2024token , or training with length-rewards RL team2025kimi ; luo2025o1 ; aggarwal2025l1 ; arora2025training ; yeo2025demystifying ; shen2025dast ; qu2025optimizing ; cui2025stepwise . Recently, ConCISE qiao2025conciseconfidenceguidedcompressionstepbystep constructs concise CoT data by inserting prompt tokens and employing early-exit during inference, then enhances the model’s reasoning conciseness through DPO rafailov2024directpreferenceoptimizationlanguage ; meng2024simposimplepreferenceoptimization . arora2025traininglanguagemodelsreason and yi2025shorterbetterguidingreasoningmodels upweight rewards for the shorter CoTs, having correct answers, in parallel-sampled reasoning chains. However, they inherit GRPO’s parallel sampling paradigm that neglects the reward function’s attention to the intermediate serial reasoning processes.
5 Conclusion
This paper proposes Serial-Group Decaying-Reward Policy Optimization, which innovatively leverages rule-based outcome rewards to regulate intermediate reasoning processes, reinforcing the model to generate high-quality thoughts earlier and exit early when the generated thought is sufficient. Experiments demonstrate that S-GRPO consistently achieves the best efficiency-accuracy synergistic improvement, significantly outperforming existing efficient reasoning methods across five benchmark datasets and four reasoning models. We recommend deploying S-GRPO as the final optimization stage in post-training pipelines.
Appendix A An example of Training Data
Here is an example that shows the training data truncation and decaying reward assignment.
Figure 5: An example shows the training data truncation and decaying reward assignment.